# Neural Representations of Dynamic Visual Stimuli

### v1

@@ 4.3 Identifying Brain Regions Tuned to Dynamic Features
Having demonstrated the ability to predict motion from fMRI brain activity, we now aim to identify brain regions selective for dynamic visual features. To do so, we test various vision encoding models to determine which best predict the fMRI brain activity of subjects watching dynamic visual stimuli. Specifically, we extract visual features using these models and construct voxel-wise encoding models using ridge regression [46, 47, 48, 49]. These voxel-wise encoding models predict the fMRI brain activity of subjects watching dynamic visual stimuli based on the extracted features.
@@
Importantly, we test models spanning multiple visual properties, such as image versus video and semantic versus visual similarity, to help identify the brain’s functional properties. By analyzing the differences in encoding capabilities, we can identify brain regions more selective to dynamic visual features.
@@
We find that the best performing video encoding model, VideoMAE Large, outperforms the best static image encoding models at predicting fMRI brain activity (Subject 1: , Subject 2: , Subject 3: , ). This suggests that fMRI data during video viewing contain dynamic information which is not captured in features from static images.
@@
Within the image encoding models, we find that CLIP ConvNeXt predicts fMRI brain activity best across all three subjects (Subject 1: , Subject 2: , Subject 3: , ), corroborating prior results [50] that guidance from semantics is useful for predicting brain responses to visual stimuli. We also find that incorporating temporal information by averaging embeddings over multiple frames from encoding models that are not trained to be aligned over time does not improve performance.
@@
We also find that the best performing embodied AI vision encoding models, VC-1, outperforms the best static image encoding models at predicts fMRI brain activity (Subject 1: , Subject 2: , Subject 3: , ). Interestingly, VC-1 has been found to be better performing than the other embodied AI vision encoding models on average across a series of tasks [3]. This suggests a possible avenue of ranking embodied AI vision encoding models on predicting fMRI brain activity.
These results suggest that models which explicitly model temporal relationships are better at capturing neural responses to dynamic visual stimuli.
Interpreting Feature Performance.
@@
We plot voxel-wise encoding performance of the best performing model, VideoMAE, on an inflated cortical map to interpret the voxel-wise selectivity for dynamic visual stimuli features in Figure 7. Please refer to the caption for more detailed analysis. We identify fMRI brain voxels that are motion selective by taking the voxel-wise encoding performance difference between the best encoding video model, VideoMAE Large, and the best encoding static image model, CLIP ConvNeXt, illustrated in Figure 8. We find results that suggest encoding visual stimuli with VideoMAE better predicts fMRI brain activity across visual cortex and in somatosensory cortex. Additionally, we show the voxel-wise encoding performance of all models on inflated cortical maps in the supplementary information.
@@ 5 Conclusion
Limitations and Future Work.
Our methods successfully decode motion of salient objects from fMRI brain activity on the Dynamic Vision Dataset [45]. However, the generalizability of our conclusions to different dataset distributions remains unclear. Future datasets with fine-grained motion can help determine the limits of motion decodability from fMRI brain activity. Furthermore, our subject-specific training approach could be improved by exploring cross-subject alignment methods, as has been shown to improve decoding performance for static stimuli [54, 55].
Conclusion and Impact.
To summarize, in this paper we propose to study neural representations of motion by disentangling the modeling of static image features and motion. We show that brain activity can be used to predict visual motion, demonstrate that our predicted motion can animate static images to reconstruct the observed dynamic visual stimuli, and identify regions in the brain that are tuned for dynamic stimuli. Our results lay the groundwork for better understanding how the brain represents visual motion, and may help researchers construct better brain-machine interfaces.
@@
Neural Representations of Dynamic Visual Stimuli
Appendix A HCP Parcellation Map
Here we show the parcellation (labeling) of the brain Figure 9.
Appendix B Additional Dataset Information
We use the Dynamic Vision Dataset collected by [45]. They release the dataset under a CC0 1.0 universal license. The url for the dataset can be found at https://purr.purdue.edu/publications/2805/1.
Appendix C Encoding Models Voxel-wise Prediction Performance on Inflated Cortex
Here we show the voxel-wise fMRI prediction performance, quantified as the Pearson correlation (r) between measured and predicted responses, for the remaining visual encoding models in alphabetical order.
Note that all following plots of encoding accuracy on inflated cortical maps are on the same scale for ease of comparison.

### v2

@@ 4.3 Identifying Brain Regions Tuned to Dynamic Features
Having demonstrated the ability to predict motion from fMRI brain activity, we identified the brain regions selective for dynamic visual features. To do so, we test various vision encoding models to determine which best predict the fMRI brain activity of participants viewing dynamic visual stimuli. Specifically, we extract visual features using these models and construct voxel-wise encoding models using ridge regression [88, 13]. These voxel-wise encoding models predict the fMRI brain activity of participants watching dynamic visual stimuli based on the extracted features.
@@
Importantly, we test models trained under different paradigms, such as those trained on static images or video input, and models that use visual or semantic contrastive losses. By analyzing the differences in brain prediction performance, we can identify brain regions more selective to dynamic visual features.
@@
We find that the best performing video encoding model, VideoMAE Large, outperforms the best static image encoding models at predicting fMRI brain activity (Pearson ’s for individual participants, S1: 0.285, S2: 0.314, S3: 0.324, all by a one-sample -test). This suggests that fMRI data during video viewing contain dynamic information which is not captured in features from static images.
@@
Within image encoding models, we find that CLIP ConvNeXt predicts fMRI brain activity best across all participants (Pearson ’s for individual participants, S1: , S2: , S3: , all ), corroborating prior results [8] that guidance from semantics is useful for visual brain prediction. Incorporating temporal information by averaging embeddings over multiple frames from encoding models that are not trained to be aligned over time does not improve performance.
@@
The best performing embodied AI vision encoding model, VC-1, outperforms the best static image encoding models at predicting brain activity (Pearson ’s for individual participants, S1: , S2: , S3: , all ). Interestingly, on average, VC-1 outperforms other embodied AI vision encoding models across a series of tasks [12]. This suggests a possible avenue for ranking embodied AI vision encoding models on predicting fMRI brain activity. More generally, models which explicitly model temporal relationships are better at capturing neural responses to dynamic visual stimuli. We provide additional controls and quantitative results in the supplemental.
@@
Interpreting Feature Performance. We plot voxel-wise encoding performance of the best performing model, VideoMAE [1], on an inflated cortical map to interpret the voxel-wise selectivity for dynamic visual stimuli features in Figure 8. Please refer to the caption for more detailed analysis. We identify fMRI brain voxels that are motion selective by taking the voxel-wise encoding performance difference between the best encoding video model, VideoMAE Large, and the best encoding static image model, CLIP ConvNeXt, illustrated in Figure 9. We find results that suggest encoding visual stimuli with VideoMAE better predicts fMRI brain activity across visual cortex and in somatosensory cortex. Voxel-wise encoding performance for each model is shown on inflated cortical maps in the supplementary information.
@@ 5 Conclusion
Limitations and Future Work. Our methods successfully decode motion of salient objects from fMRI brain activity on the Dynamic Vision Dataset [87]. However, the generalizability of our conclusions to different dataset distributions remains unclear. Future datasets with fine-grained motion can help determine the limits of motion decodability from fMRI brain activity. Furthermore, our subject-specific training approach could be improved by exploring cross-subject alignment methods, as has been shown to improve decoding performance for static stimuli [92, 93].
Conclusion and Impact. To summarize, in this paper we propose to study neural representations of motion by disentangling the modeling of static image features and motion. We show that brain activity can be used to predict visual motion, demonstrate that our predicted motion can animate static images to reconstruct the observed dynamic visual stimuli, and identify regions in the brain that are tuned for dynamic stimuli. Our results lay the groundwork for better understanding how the brain represents visual motion, and may help researchers construct better brain-machine interfaces.
@@
Reanimating Images using Neural Representations of Dynamic Stimuli
Appendix A Supplemental
Sections
A.1 HCP Parcellation Map
A.2 Additional Dataset Information
We use the Dynamic Vision Dataset collected by Wen et al. [87]. The dataset is released under a CC0 1.0 universal license and can be found at https://purr.purdue.edu/publications/2805/1.
A.3 Decoding Models: Visualizations
Our website https://sites.google.com/view/neural-dynamics/home has multiple reanimation examples as videos. For each example, we show the ground truth video, the reanimated video using the flow predicted from the ground truth initial frame, and the reanimated video using the flow predicted from the initial frame generated by MindVideo [3]. The MindVideo initial frame represents the diffusion image as predicted by MindVideo from fMRI data. As such, this frame is often quite different in content from the ground truth, although the objects in each are in similar locations. However, the motion predicted by our model is still consistent with ground truth motion, for example, in the first video the jellyfish retracts backwards at the end clip. The same backwards motion can be seen in the reanimated videos for with both the ground truth initial frame and the MindVideo generated initial frame (which happens to be a boat). This establishes that the motion visualized using DragNUWA [4] is not derived solely from the diffusion model, but rather, incorporates our predicted motion.
In addition, we display example generated video frames below. Within each figure the first row is the ground truth frame, the second row is our reanimated video using the flow predicted from the ground truth initial frame, and the final row is the reanimated video using the flow predicted from the initial frame generated by MindVideo.
A.4 Encoding Models: Controls and Baselines
| VideoMAE Models (Pearson ) | ||
| Model | Dataset | Encoding Performance |
| VideoMAE Base | K-400 | 0.2964 |
| VideoMAE Large | K-400 | 0.3248 |
| VideoMAE Base | SSV2 | 0.3137 |
| Hiera Models (Pearson ) | ||
| Model | Dataset | Encoding Performance |
| Hiera Base Plus | K-400 | 0.2629 |
| Hiera Base Plus | IN-1K | 0.2198 |
| Hiera Huge | K-400 | 0.2532 |
| Hiera Huge | IN-1K | 0.2407 |
A.5 Encoding Models: Voxel-wise Prediction Performance on Inflated Cortical Maps of the Brain
Inflated brain maps showing voxel-wise fMRI prediction performance, quantified as the Pearson correlation () between measured and predicted responses, for all visual encoding models not shown in the main text. Maps for each model are shown in alphabetical order. ‘AF’ in the model name denotes a model using Average Frames. All maps are shown using the same scale for ease of comparison between models.